Izmir A.S. issued convertible bonds at their face value of 109,000 lira on December 31, 2017. The bonds have a 10-year life with interest of 10 percent payable annually. At the date of issue, the prevailing interest rate for similar debt without a conversion option was 12 percent.

Assume that a foreign company using IFRS is owned by a company using U.S. GAAP. Thus, IFRS balances must be converted to U.S. GAAP to prepare consolidated financial statements. Ignore income taxes.

Required:

  1. a. Prepare journal entries for this compound financial instrument for the year ending December 31, 2017, under (1) IFRS and (2) U.S. GAAP.

  2. b. Prepare the entry(ies) that the U.S. parent would make on the December 31, 2017, conversion worksheet to convert IFRS balances to U.S. GAAP.
Answer #1

Answer a.

Part 1 - Under IFRS

Account title Debit Credit
Cash 109000
Bonds Payable 96682
Additional paid in capital - convertible bonds 12318

pv value of bonds payable: [109000 *0.3219732] + [10900 * 5.6502] = 35095 + 61587 = 96682

Note: The above amount can also be 96685 and 12315, if only upto four decimal places are taken.

Leave Comment, in case of any doubt.

Part 2 - Under U.S. GAAP

Account title Debit Credit
Cash 109000
Bonds payable 109000

Answer b.

Conversion from IRFS TO GAAP

Account title Debit Credit
Additional paid in capital - convertible bonds 12318
Bonds payable 12318